Circular No. 6267

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1995ao AND 1995ap IN ANONYMOUS GALAXIES

R. P. Kirshner, Center for Astrophysics (CfA), reports the discovery of two supernovae on behalf of the High-Z Supernova Search Team. The first object (SN 1995ao) is located at α = 2^{h}57^{m}30^{s}.70, δ = −1^{o}41′19′′.8 (equinox 2000.0); a nearby foreground star has position end figures 32^{s}.15, 40′43′′.5. The second object (SN 1995ap) is located at α = 3^{h}12^{m}28^{s}.13, δ = +0^{o}41′43′′.4; a nearby star has position end figures 27^{s}.23, 55′′.7. Both supernovae were discovered on frames taken by M. Phillips (Cerro Tololo) and B. Schmidt (Mount Stromlo and Siding Spring Observatories) with the Cerro Tololo 4-m telescope on Nov. 18 UT at R ∼ 21.5. SN 1995ao was not present to R = 22.5 on a similar frame taken on Oct. 28. SN 1995ap was present at R ∼ 22 on a frame taken on Oct. 28. Spectra obtained at the Multiple Mirror Telescope (MMT) by R. Kirshner and R. Jayawardhana (CfA) on Nov. 23, and at the Keck telescope by A. V. Filippenko, S. Perlmutter, A. J. Barth, and I. Hook (University of California, Berkeley), show that SN 1995ao is a type-Ia event near maximum. The redshift of the supernova spectrum from the MMT and from Keck agrees with that of the host galaxy, derived from emission lines in the Keck spectrum, at z = 0.24. Spectra of SN 1995ap from both the MMT and Keck telescopes indicate that it is also of type Ia near maximum. The redshift of the supernova, derived from its spectrum, is ∼ 0.3. Finding charts of these two objects are available via e-mail from brian@mso.anu.edu.au. Photometric observations of these objects in V, R, and I would be valuable.

GRS 1915+105 S. S. Eikenberry and G. G. Fazio, CfA, report on near-infrared observa-

tions of GRS 1915+105 during its recent period of hard x-ray/radio flaring (IAUC 6266), using the Kitt Peak 2.1-m telescope (+ COB infrared-array camera): “The following magnitudes were obtained under photometric conditions on Oct. 17.13 UT: J = 17.81 ± 0.17, H = 15.16 ± 0.04, K = 13.43 ± 0.03. Under nonphotometric conditions, we obtained the following magnitudes on Oct. 16.09 using differential photometry with other stars in the field: H = 14.94 ± 0.09, K = 13.51 ± 0.05. The magnitudes and colors closely match those observed by Mirabel et al. (1994, A.Ap. 282, L17) during similar hard x-ray/radio flaring activity on 1993 June 5. Our data do not reveal any significant night-to-night variations, in contrast to the night-to-night variations of ∼ 1 magnitude observed by Mirabel et al.”

1995 November 28 Daniel W. E. Green